# Spica: Retrieving Scenarios for Pluralistic In-Context Alignment

### v1

@@ 5.3 Generalizing to Non-Discrete Preferences
While this work explores in-context learning approaches to value alignment, the models we use as a source to build aligned models from also come with their own existing safeguards. In the case of gpt-4o, this can be especially challenging, as the model itself is not open source and only being able to interact with such models through a proprietary API likely leaves the possibility of external safeguards. These safeguards are often implemented as additional models that may monitor the inputs and outputs around the main LLM, jumping in to intervene should it identify inappropriate inputs being supplied or outputs being generated. This means our ability to affect the outputs of such models may be limited in ways that cannot be addressed by modifying the prompts.
Strategy Adherence
In our study, we use a set of 5 response strategies to approximate a diverse set of responses for each prompt. While there is evidence from prior work that human preferences tend to align towards high-level strategies Cheong et al. (2024), generating responses following fixed strategies may not always be reliable as responses may not always adhere to the strategies (either due to model safeguards or relevance of the strategy to an input prompt). To control for the effects of this, during our annotations of the scenario bank, we asked annotators for input on both concrete responses and high-level instructions and only used the corresponding rating data when testing prompting strategies based on instructions versus examples. Still, this may be insufficient to address the resulting reduction in variation of the response space on some prompts.
@@
In our experiments, we focused primarily on a small-scale proof-of-concept alignment task targeted towards a US population. As a result, we were only able to examine the outcomes of alignment over one source of input prompts (PRISM) and several demographically-constructed groups based on US participants. While in this setup, we did observe differences between alignment mechanisms and also differences between goals (pluralistic alignment at a group level versus alignment towards the entire population), it is likely that given different constructions of the groups, there may be different takeaways.
@@
The AI alignment problem itself has many ethical implications, and these considerations also extend to both implications of the the design of Spica, and our choices during our evaluation of it.
First, our experiments are intended to demonstrate a proof-of-concept setting where different groups are likely to have significant divergent values, as a result of this consideration and practicalities surrounding ease of recruitment, we we opted to extrinsically define “groups” based on divisive demographic features within a US-based participant pool. However, this should not be interpreted as an endorsement for using politics and religion as a way to conduct pluralistic alignment—many other factors like culture, community, and identity could provide better delineation between different groups with lower risks around introducing additional social fragmentation. Given this, we also caution against using results in this work to make inferences about the broader population groups we tested with, as we didn’t make additional efforts to ensure our participants are representative samples within these groups.
@@ A.1 Results: Group-level Breakdown of the Retrieval Loss
We present a group-by-group breakdown of the retrieval loss in Figure 7. Interestingly, we find that the groups indicting higher affinity to religion (-rel) tended to see a more marked difference in retrieval quality. This seems to be the result of these groups having more preferences over responses that are not as dependent on the specific prompt and instead apply to a wide variety of topics. For gpt-4o, the P-I and C-R conditions consistently produced positive alignment outcomes.
@@ A.3 Results: Group-level Breakdown of End-to-End Evaluation
We present a group-by-group breakdown of the final end-to-end evaluation in Figure 9. For gpt-4o, we found Spica with contrastive examples to be the provide the most consistent alignment across groups, being preferred over the control response, but not always the most preferred response across the alignment conditions. Baseline retrieval seemed to be effective in alignment for dem-identifying groups but produced the opposite outcome for rep-identifying ones.
A.4 Results: Qualitative Analysis of the Weights Learned
@@
Finally, we qualitatively look at the weights learned for various groups across each model. Here we observe that, weights produced after learning from response types preferences and response example preferences end up relatively similar to each other. We also note that similarity scores (in this case cosine similarity) receives a comparatively lower absolute weight compared to the other metrics, however, this is as expected, as similarity scores tend to span a different range of values than preference level metrics. We also observe that between the two new metrics, stability is the most important for the all experiment, matching the notion that in a collective alignment setting, using examples that are closer to universal values tends to be more ideal, while at a group-level there is no such pattern. Finally, for the -nrel groups we observed cases where similarity was assigned a positive weight, implying that examples immediately closer to the query were actually often less desirable, possibly a reflection of non-religious groups finding subject matter around different religious topics less similar to each other than religious identifying groups. Beyond this, though, the weights seem overall unsurprising, with no other significant patterns of note.
A.5 Human Annotation Materials
In this section we document the instructions and materials used for our human annotation and evaluation tasks.
A.5.1 Instructions
@@ A.5.4 Questions for Evaluation
A.6 Prompts and Instructions
A.6.1 Response Strategies
With scenario banks, user preferences can be assessed over both strategies that cover a high-level style for responding and concrete responses produced by a model. In this appendix section, we provide the categories we used to delineate different high-level response strategies (adapted from findings in Cheong et al. (2024)) and also provide a set of associated prompts that we used for each strategy to generate concrete model responses.
Strategy Description Prompt
@@ A.6.1 Response Strategies
Table 1: Prompts used to populate the responses for scenarios in the scenario bank based on 5 general strategies. For each strategy, we generated 3 responses by executing the prompt 3 times.
A.6.2 ICL Alignment Prompts
@@
This prompt by itself was used to generate the control response used during evaluations, and is meant to represent a setup for a default alignment.
@@
In each case the APPROPRIATENESS label uses the rating description (Table 1) that most closest matches the appropriateness of the best (highest rated) and worst (lowest rated) response or strategy.

### v2

@@ 5 Conclusions and Discussion
While this work explores in-context learning approaches to value alignment, the models we use as a source to build aligned models from also come with their own existing safeguards, particularly for closed-source models like gpt-4o. This means our ability to affect the outputs of such models may be limited in ways that cannot be addressed by prompt-based steering.
Adherence to Response Classes
In our study, we use a set of 5 response classes (and associated prompts) to approximate a diverse span of possible responses for each prompt. While there is evidence from prior work that human preferences tend to align towards these high-level classes of responses Cheong et al. (2024), generating responses following fixed strategies may not always be reliable, as actual responses may not always adhere to the strategies for each class (either due to model safeguards or relevance of the strategy to an input prompt). To control for the effects of this, during our annotations of the scenario bank, we asked annotators for input on both concrete responses and high-level instructions and only used the corresponding rating data when testing prompting strategies based on instructions versus examples. Still, this may be insufficient to address the resulting reduction in variation of the response space on some prompts. Future work can explore alternative categories that do not constrain the response space in the same way.
@@
In our experiments, we focused primarily on a small-scale proof-of-concept alignment task targeted towards a US population. As a result, we were only able to examine the outcomes of alignment over one source of input prompts (PRISM) and several demographically-constructed groups based on US participants. While in this setup, we observed differences between alignment mechanisms and goals (e.g., group-level pluralistic alignment vs. population-wide alignment), different group configurations could yield different takeaways.
@@
The AI alignment problem itself has many ethical implications, and these considerations also extend to both implications of the design of Spica, and our choices during our evaluation of it.
First, our experiments are intended to demonstrate a proof-of-concept setting where different groups are likely to have significant divergent values. As a result of this consideration and practicalities surrounding ease of recruitment, we we opted to extrinsically define “groups” based on divisive demographic features within a US-based participant pool. However, this should not be interpreted as an endorsement for using politics and religion as a way to conduct pluralistic alignment—many other factors like culture, community, and identity could provide better delineation between different groups with lower risks around introducing additional social fragmentation. Given this, we also caution against using results in this work to make inferences about the broader population groups we tested with, as we didn’t make additional efforts to ensure our participants are representative samples within these groups.
@@ A.1 Results: Group-level Breakdown of the Retrieval Loss
indicating
We present a group-by-group breakdown of the retrieval loss in Figure 6. Interestingly, we find that the groups indicating higher affinity to religion (-rel) tended to see a more marked difference in retrieval quality. This seems to be the result of these groups having more preferences over responses that are not as dependent on the specific prompt and instead apply to a wide variety of topics. For gpt-4o, the P-I and C-R conditions consistently produced positive alignment outcomes.
@@ A.3 Results: Group-level Breakdown of End-to-End Evaluation
We present a group-by-group breakdown of the final end-to-end evaluation in Figure 8. For gpt-4o, we found Spica with contrastive examples to provide the most consistent alignment across groups, being preferred over the control response, but not always the most preferred response across the alignment conditions. Baseline retrieval was observed as effective in alignment for dem-identifying groups but produced the opposite outcome for rep-identifying ones.
A.4 Results: Qualitative Analysis of Learned Weights
@@
Finally, we qualitatively look at the weights learned for various groups for each model. Here we observe that weights produced after learning from response types preferences and response example preferences end up relatively similar to each other. We also note that similarity scores (in this case cosine similarity) receive a comparatively lower absolute weight compared to the other metrics. However, this is as expected, as similarity scores tend to span a different range of values than preference level metrics. We also observe that between the two new metrics, stability is the most important for the all experiment, matching the notion that in a collective alignment setting, using examples that are closer to universal values tends to be more ideal, while at the group level there is no such pattern. Finally, for the -nrel groups we observed cases where similarity was assigned a positive weight, implying that examples immediately closer to the query were actually often less desirable, possibly a reflection of non-religious groups finding subject matter around different religious topics less similar to each other than religious identifying groups. However, beyond this, the weights seem generally unsurprising, with no other significant patterns of note.
A.5 Results: Pluralistic versus Collective Alignment
We observe (Figure 10) that, unlike in the setting with separate groups, optimal prompt strategies now significantly favor instructions (P-I and C-I) on gpt-4o, likely due to none of the examples being good candidates to represent collective values. On the end-to-end evaluation of the test set queries, also perhaps unsurprisingly, group-informed retrieval metrics from Spica no longer seem to provide any significant benefit, even slightly under-performing baseline retrieval. We attribute this to the fact consistent norms are unlikely in the collective group, leaving little benefit to using group-informed retrieval metrics, coupled with Spica-R no longer reflecting an effective prompting setup in this setting. In fact, for the collective case, the ICL prompt style becomes the most important factor, with gpt-4o favoring instructions and llama3-8b now favoring example responses (baseline and Spica-R).
A.6 Human Annotation Materials
In this section, we document the instructions and materials used for our human annotation and evaluation tasks.
A.6.1 Instructions
@@ A.6.4 Questions for Evaluation
A.7 Prompts and Instructions
A.7.1 Response Classes
With scenario banks, user preferences can be assessed over both strategies that cover a high-level style for responding and concrete responses produced by a model. In this appendix section, we provide the categories we used to delineate different high-level response classes and associated strategies (adapted from findings in Cheong et al. (2024)) and also provide a set of associated prompts that we used to generate concrete model responses.
Response Class Description Prompt
@@ A.7.1 Response Classes
Table 1: Prompts used to populate the responses for scenarios in the scenario bank based on 5 general classes of responses. For each class of response, we generated 3 responses by executing the prompt 3 times.
A.7.2 ICL Alignment Prompts
@@
This prompt by itself was used to generate the control response used during evaluations and is meant to represent a setup for a default alignment.
@@
In each case the APPROPRIATENESS label uses the rating description (Table 1) that most closely matches the appropriateness of the best (highest rated) and worst (lowest rated) response or strategy.